OpenStudy (anonymous):

What is the domain and range of the quadratic equation y = –2x^2 – 32x – 126?

OpenStudy (experimentx):

the domain is set of real numbers

OpenStudy (anonymous):

Do you know the range?

OpenStudy (experimentx):

for range let's see ... your equation is parabolic equation facing downwards, find the vertex first ... vertex is going to be -32/4 = -8 ... so range is set of real numbers less or equal to -8

OpenStudy (anonymous):

What about this one? What are the x– and y–intercepts of the quadratic equation y = 2x^2 – 8x + 6?

OpenStudy (experimentx):

the domain is same ... range is set of real numbers equal or greater than -b/2a = 2

OpenStudy (anonymous):

It's asking about the x and y intercepts not domain and range lol :P

OpenStudy (experimentx):

you will have 2 x intercepts, and 1 y intercept y intercept is 6, for x-intercept put y=0, you have to solve 2x^2 – 8x + 6 = 0

OpenStudy (anonymous):

Thank you SO much! I finally understand :D